OpenStudy (anonymous):

The number of sofas a factory produces varies directly with the number of hours the machinery is operational. Suppose the factory can produce 455 sofas in 48 hours. What is an equation that relates the number of sofas produced, n, with the amount of time, t, in hours? What is the graph of your equation? A. n = 9.48t B. n = 0.11t C. t = 4.74n D. t = 9.48n

OpenStudy (anonymous):

i know it is not (D)

OpenStudy (ghazi):

455 sofas in 48 hours 1 sofa in 48/455 hours= 0.105 hours

OpenStudy (ghazi):

or in 48 hours 455 sofa in 1 hour number of sofa= 455/48=9.47 sofas

OpenStudy (ghazi):

therefore t=9.48 n

OpenStudy (anonymous):

so it is (A)

OpenStudy (ghazi):

it's D

OpenStudy (anonymous):

i got that wrong when i put (D) =(

OpenStudy (anonymous):

They probably used t for time and n for number of sofas.

OpenStudy (anonymous):

i think (B)

OpenStudy (ghazi):

then i guess ..may be there answer is wrong ...because whole explanation is in front of you :)

OpenStudy (ghazi):

no it can't be B or C

OpenStudy (anonymous):

then i will go with (A)

OpenStudy (anonymous):

n=kt 455=k(48) k=455/48 n=(455/48)t n=9.48t

OpenStudy (anonymous):

thanks =)

OpenStudy (anonymous):

You're welcome.